# The discrete periodic Pitman transform: invariances, braid relations, and Burke properties

### v1

@@ A.1. Basic algebraic properties
We adapted the definition of the periodic Pitman transform from the work of [CGS24]
$\mathfrak{T}_{i}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})=X_{1,i}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,[i,j]}-X_{1,[i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,[i+1,j]-X_{1,[i+1,j]}}}}\Biggr{)}$
$\mathfrak{D}_{i}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})=X_{2,i}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,(i,j]}-X_{1,(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,(i-1,j]}-X_{1,(i-1,j]}}}\Biggr{)}$
In [CGS24], these maps were denoted as $T^{N,2}$ and $D^{N,2}$, respectively. It is important to note that this is equivalent to our definition of the map up to a shift:
@@
where $\tau_{k}:\mathbb{R}^{\mathbb{Z}_{N}}\rightarrow\mathbb{R}^{\mathbb{Z}_{N}}$ is the shift map: $(\tau_{k}X)_{i}=X_{i+k}$ for all $i\in\mathbb{Z}_{N}$.
@@
$\frac{1}{\beta}\log\Bigl{(}\sum\limits_{i=0}^{k}\exp(\beta E_{i})\Bigr{)}$ converges to $\max\limits_{i}E_{i}$, uniformly on the values of $E$ (but not k) as $\beta\to\infty$.
@@
$=\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}$
@@
$\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}$ $\leq\frac{1}{\beta}\log\Biggl{(}\frac{k\max\limits_{i}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}=\frac{1}{\beta}\log(k)$
@@
$\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}$ $\geq\frac{1}{\beta}\log\Biggl{(}\frac{\max\limits_{i}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}=\frac{1}{\beta}\log(1)=0$
@@
The maps $\mathcal{P}$ and $\widetilde{\mathcal{P}}$ define in (1.2) and (1.11) are involutions on $\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$.
@@
    If $(\mathbf{U}_{1},\mathbf{U}_{2})=\mathcal{P}(\mathbf{X}_{1},\mathbf{X}_{2})$, then $\mathcal{P}(\tau_{1}\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2})=(\tau_{1}\mathbf{U}_{1},\tau_{1}\mathbf{U}_{2})$
@@
    If $(\mathbf{U}_{1},\mathbf{U}_{2})=\widetilde{\mathcal{P}}(\mathbf{X}_{1},\mathbf{X}_{2})$, then $\widetilde{\mathcal{P}}(\tau_{1}\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2})=(\tau_{1}\mathbf{U}_{1},\tau_{1}\mathbf{U}_{2})$
@@
$D(\mathbf{X}_{1},D(\mathbf{X}_{2},\mathbf{X}_{3}))=D\bigl{(}D(\mathbf{X}_{1},\mathbf{X}_{2}),D(T(\mathbf{X}_{1},\mathbf{X}_{2}),\mathbf{X}_{3})$

### v2

@@ A.1. Basic algebraic properties
In the present work, we use a different formulation of the maps $D$ and $T$ than those introduced in [CGS24]; there, the relevant mappings were defined as
$\mathfrak{T}_{i}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})$ $=X_{1,i}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,[i,j]}-X_{1,[i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,[i+1,j]-X_{1,[i+1,j]}}}}\Biggr{)},\quad\text{and}$
$\mathfrak{D}_{i}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})$ $=X_{2,i}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,(i,j]}-X_{1,(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,(i-1,j]}-X_{1,(i-1,j]}}}\Biggr{)}.$
In [CGS24], these maps were denoted as $T^{N,2}$ and $D^{N,2}$, respectively. It is important to note that this is equivalent to our definition of the maps defined in (1.2) up to a shift:
@@
where, for $k\in\mathbb{Z}_{N}$, $\tau_{k}:\mathbb{R}^{\mathbb{Z}_{N}}\rightarrow\mathbb{R}^{\mathbb{Z}_{N}}$ is the shift map: $(\tau_{k}X)_{i}=X_{i+k}$ for all $i\in\mathbb{Z}_{N}$.
@@
As $\beta\to\infty$, $\frac{1}{\beta}\log\Bigl{(}\sum\limits_{i=0}^{k}\exp(\beta E_{i})\Bigr{)}$ converges to $\max\limits_{i}E_{i}$, uniformly as a function of the $E_{i}$ (but not as a function of $k$)
@@
$=\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}.$
@@
$\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}$ $\leq\frac{1}{\beta}\log\Biggl{(}\frac{k\max\limits_{i}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}=\frac{1}{\beta}\log(k).$
@@
$\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}$ $\geq\frac{1}{\beta}\log\Biggl{(}\frac{\max\limits_{i}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}=\frac{1}{\beta}\log(1)=0.$
@@
The maps $\mathcal{P}$ and $\widetilde{\mathcal{P}}$ defined in (1.2) and (1.13) are involutions on $\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$.
@@
    If $(\mathbf{U}_{1},\mathbf{U}_{2})=\mathcal{P}(\mathbf{X}_{1},\mathbf{X}_{2})$, then $\mathcal{P}(\tau_{1}\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2})=(\tau_{1}\mathbf{U}_{1},\tau_{1}\mathbf{U}_{2})$.
@@
    If $(\mathbf{U}_{1},\mathbf{U}_{2})=\widetilde{\mathcal{P}}(\mathbf{X}_{1},\mathbf{X}_{2})$, then $\widetilde{\mathcal{P}}(\tau_{1}\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2})=(\tau_{1}\mathbf{U}_{1},\tau_{1}\mathbf{U}_{2})$.
@@
$D(\mathbf{X}_{1},D(\mathbf{X}_{2},\mathbf{X}_{3}))=D\bigl{(}D(\mathbf{X}_{1},\mathbf{X}_{2}),D(T(\mathbf{X}_{1},\mathbf{X}_{2}),\mathbf{X}_{3}).$
@@
Appendix B Proofs of the zero-temperature analogues in Theorems 1.10-1.12
Lemma B.1.
For any $k\in\mathbb{N}$ and any pair of $k$-tuples of ordered pairs $U=\bigl{(}(a_{i},b_{i})\bigr{)}_{i\in\llbracket 1,k\rrbracket}$, $V=\bigl{(}(c_{i},d_{i})\bigr{)}_{i\in\llbracket 1,k\rrbracket}$, as $\beta\to\infty$, $\frac{1}{\beta}\log\bigl{(}Z^{\beta,\mathbf{W}}(V\mid U)\bigr{)}$ converges uniformly (as a function of $\mathbf{W}$) to $G^{\mathbf{W}}(V\mid U)$.
Proof.
This follows as a direct consequence of Lemma A.1 since there are finitely many multi-paths between any two sets of points. ∎
Proof of Theorem 1.10..
$\quad\,\mathcal{P}_{k}^{\beta}(\ldots,\mathbf{W}_{k-1},\mathbf{W}_{k},\mathbf{W}_{k+1},\mathbf{W}_{k+2},\ldots)$ (B.1)
$Z^{\beta,\mathbf{W}}(V\mid U)=Z^{\beta\mathbf{W}}(V\mid U)=Z^{\mathcal{P}_{k_{1}}(\beta\mathbf{W})}(V\mid U)=Z^{\beta(\frac{1}{\beta}\mathcal{P}_{k_{1}}(\beta\mathbf{W}))}(V\mid U)=Z^{\beta,\mathcal{P}_{k_{1}}^{\beta}(\mathbf{W})}(V\mid U).$ (B.2)
By Lemma B, we know that $\frac{1}{\beta}\log\bigl{(}Z^{\beta,\mathbf{W}}(V\mid U)\bigr{)}$ converges to $G^{\mathbf{W}}(V\mid U)$ uniformly on the weights, so $G^{\mathbf{W}}(V\mid U)$ is preserved under the limiting operator $\lim\limits_{\beta\to\infty}\mathcal{P}_{k_{n}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}$. We claim that each $\mathcal{P}_{k}^{\beta}$ converges to $\widetilde{\mathcal{P}}_{k}$ uniformly as a function of the weights, which implies by induction that
$G^{\mathbf{W}}(V\mid U)=G^{\lim\limits_{\beta\to\infty}\mathcal{P}_{k_{m}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}\mathbf{W}}(V\mid U)=G^{\widetilde{\mathcal{P}}_{k_{m}}\ldots\widetilde{\mathcal{P}}_{k_{1}}\mathbf{W}}(V\mid U),$ (B.3)
$=\Bigl{(}\frac{1}{\beta}\log\bigl{(}\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i,j]}}\bigr{)}-\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\Bigr{)}-\Bigl{(}\frac{1}{\beta}\log\bigl{(}\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i-1,j]}}\bigr{)}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}\Bigr{)}.$
We next show the braid relations pass to zero-temperature.
Proof of Theorem 1.11..
Defining $\mathcal{P}_{k}^{\beta}$ as in (B.1), we see that $\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}=\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}$, as an immediate consequence of Theorem 1.6. As noted in the proof of Theorem 1.10, $\mathcal{P}_{k}^{\beta}$ converges to $\widetilde{\mathcal{P}}_{k}$ uniformly. This implies the braid relation for $\widetilde{\mathcal{P}}_{k}$, since we have
To prove Theorem 1.12, we first need analogous Burke properties in the geometric and exponential cases.
Proposition B.2 (Geometric Burke property).
$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})).$
Proof.
Proposition B.3 (Exponential Burke property).
Let $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{Z}_{\geq 0}^{\mathbb{Z}_{N}}$ be independent sequences of independent random variables such that $X_{1,i}\sim\text{Exp}(a_{1}+b_{i})$ and $X_{2,i}\sim\text{Exp}(a_{2}+b_{i})$ for all $i\in\mathbb{Z}_{N}$ where $a_{m}+b_{i}>0$ for $m\in\{1,2\}$ and $i\in\mathbb{Z}_{N}$. Then,
$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})).$
Proof.
For $m\in\{1,2\}$ and $i\in\mathbb{Z}_{N}$, define
$X_{m,i}^{\prime}\sim\text{Geom}\Biggl{(}\left(1-\frac{a_{m}}{n}\right)\left(1-\frac{b_{i}}{n}\right)\Biggr{)}.$
$\frac{1}{n}\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})=\widetilde{T}\Bigl{(}\frac{\mathbf{X}_{1}^{\prime}}{n},\frac{\mathbf{X}_{2}^{\prime}}{n}\Bigr{)},\quad\text{and}\quad\frac{1}{n}\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})=\widetilde{D}\Bigl{(}\frac{\mathbf{X}_{1}^{\prime}}{n},\frac{\mathbf{X}_{2}^{\prime}}{n}\Bigr{)}.$
By Proposition B.2, we know that $(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}),\widetilde{D}(\mathbf{X}_{1}^{\prime},\mathbf{X}_{2}^{\prime}))$. Since $\max$, addition, and subtraction are continuous, $\widetilde{D}$ and $\widetilde{T}$ are continuous maps. We then have that
Proof of Theorem 1.12.
The proof is the same as the proof of Theorems 1.7 and 1.8: replace the use of Theorem 1.4 with Theorem 1.10 and replace the Burke property in Propositions 4.1 with the Burke property in Propositions B.2 and B.3. ∎